OpenStudy (anonymous):

use a formula to find the width w of the rectangle that has a perimeter of 24cm nd a length of 9cm

OpenStudy (anonymous):

P=2w+2L 2w=P-2L w=(P-2L)/2 w=(24-2(9))/2 w=3